# BA-SAM: Scalable Bias-Mode Attention Mask for Segment Anything Model

### 04 Jan 2024 v1

@@ Abstract
Figure 1: (a) Top: contrast between prior methods [79, 4] and BA-SAM. For large-scale datasets, previous approaches often resize images or change patch sizes to handle the issue of varying resolutions. In contrast, we propose a Scalable Bias-Mode Attention Mask (BA-SAM), which enhances SAM’s adaptability to varying image resolutions while eliminating structure modifications. Bottom (left): We introduce a generalized model that does not require any task-specific designs, yet outperforms state-of-the-art methods across four datasets. Bottom (right): With resolution variations, prior models’ performance degrades drastically. Instead, BA-SAM consistently alleviates this issue.
@@ 1 Introduction
Recently, the computer vision community [30, 66, 29, 9, 74, 26, 85, 84, 86, 83, 87, 88, 89, 31, 90, 22, 58, 50, 49] has experienced a surge in the development of various foundation models [54, 35, 21]. Notably, Meta has introduced SAM (Segment Anything Model) [38], a prompt model that has made a significant impact. SAM can segment any object in an image or video by incorporating a single visual prompt, such as a box or a point, without requiring additional training. SAM is trained on an extensive SA-1B dataset [38], consisting of over 11 million images and one billion masks. Its emergence has undeniably showcased robust generalization capabilities across diverse images and objects, paving the way for new possibilities and avenues in intelligent image analysis and understanding [76, 8, 79, 34]. Based on SAM, Some variants have been proposed, such as MobileSAM [76] and SAM-Adapter [8]. These efforts typically focus on improving SAM’s performance on specific datasets.
@@
Since SAM adopts the standard Vision Transformer [14] architecture, there are two common approaches to address the inconsistency between training and inference sizes for the ViT architecture. As depicted in Figure 1, the first approach, e.g., MSA [79] and SAM-Adapter [8], involves directly resizing all datasets to match a predefined size. Conversely, the second approach, exemplified by FlexiViT [4], entails adjusting the patch size to accommodate larger image resolutions. Nevertheless, tuning the image or patch size necessitates a complete retraining of the model, which is cost-expensive and unacceptable for deployment in the downstream tasks. Besides, it prevents leveraging the rich prior knowledge reserved in the pre-trained model of SAM. As such, our objective is to explore a solution that enhances SAM’s adaptability to datasets of varying resolutions while avoiding structural modifications to SAM.
In this paper, we introduce a novel perspective that reframes the challenge of image resolution variation as a length extrapolation problem. Specifically, as depicted in Figure 1, for images of varying sizes, we employ different token sequence lengths while keeping a consistent patch size. It has been observed that the inconsistency in token length between training and prediction is a key factor in performance degradation. This inconsistency manifests in two aspects: Firstly, changes in token length lead to variations in the magnitude of attention module values. When the dot product result becomes significantly large in magnitude, it can drive the subsequent Softmax layer into regions with minimal gradients. Consequently, the attention distribution after Softmax becomes highly concentrated, giving rise to the issue of vanishing gradients. Secondly, longer predictions rely on untrained information, such as additional position encodings. The introduction of untrained parameters brings a considerable amount of noise to the model, which, in turn, affects its performance.
@@
Our approach demonstrates efficacy in two scenarios: zero-shot and fine-tuning. Extensive evaluation on datasets from five diverse tasks are conducted, including DIS5K [53], DUTS [64], ISIC [12], COD10K [16], and COCO [45]. These datasets vary in resolution, mostly exceeding SAM’s default resolution of 1024. In the zero-shot setting, our BA-SAM alleviates the model’s performance degradation caused by expanding the inference resolution without requiring any additional training. With a few fine-tuning epochs on downstream tasks, our BA-SAM consistently achieves state-of-the-art accuracy across all datasets. Additionally, to further demonstrate BA-SAM’s generalizability, we propose a generalized model and a new benchmark, which utilize one model to attain state-of-the-art performance across all four datasets simultaneously.
@@ 2 Related Work
Visual Foundation Models. Models that are trained on broad and can be adapted to numerous downstream tasks are called “Foundation Models” [5, 67, 43]. These models, Vision-Language Models (VLM) (CLIP [54] and DALL-E [55]) combine computer vision and natural language processing to understand and generate descriptions or analyze visual content using textual and visual information. Masked Image Modeling [73, 48] (MIM) refers to a technique where parts of an image are masked during training to encourage a model to learn contextual information and complete missing regions. SAM [38] is a model designed for segmenting objects or areas in images, offering precise segmentation capabilities. In this paper, we use a variant of SAM called MobileSAM [76] as the baseline method. As our design optimizes the Transformer itself, it applies to all Transformer-based vision foundation models.
@@
Figure 2: Illustration of the proposed BA-SAM method. (a) In the original SAM, when input token sequences length varies during testing, the magnitude of the Softmax outputs changes drastically. In contrast, we propose a new scaling factor to address this issue. (b) We introduce a bias-mode attention mask, which produces increasing penalties on attention scores as the distance between query and key grows.
@@
Resolution Variation Processing. To enable models to be more adaptable to variations in resolutions, traditional methods to deal with VIT have relied on adjustments to positional embeddings [40] and patch sizes [4, 6, 24, 75, 44, 30, 39]. Patch n’ Pack [13] employed sequence packing during training to handle inputs with arbitrary resolutions and aspect ratios. All of them necessitate training from scratch, incurring substantial computational and time costs. In contrast to previous approaches, we extend the concept of ’length extrapolation’ from NLP into the context of addressing scale variations in CV. Length extrapolation refers to a model’s ability to generalize well to longer inputs than those it was trained on. In NLP, it has been successfully used, such as in ALIBI [52] and KERPLE [11], to enable models to adapt to longer sequences without significant performance degradation. Our approach seamlessly extends to two scenarios: zero-shot and finetuning, allowing us to leverage prior knowledge embedded in the SAM and significantly reduce training efforts.
@@
Parameter Efficient Tuning. There have been some pioneering works for the Parameter Efficient Tuning (PETL) of visual models, such as AdaptFormer [7] and visual prompt tuning (VPT) [36]. He et al. [28] analyzed the unified view among PETL techniques such as prefixtuning [42], Prompt-tuning [36], and adapter [7]. Our method belongs to the category of Parameter Efficient Tuning.
@@
Visual Attention Modeling. Various studies have incorporated attention mechanisms into neural network architectures designed for visual tasks [63, 68, 41, 3, 33, 93]. These mechanisms are employed in a channel-wise manner to capture cross-feature information [66, 10, 80]. They are also used for selecting paths in different branches of a network [61], or a combination of both strategies [78]. For instance, squeeze-and-excite network [32] features an attention-like module for modeling channel-wise relationships within layer features. Li et al. [41] utilize the attention mechanism to adapt the receptive field of neurons between network branches. The advent of transformers has led to hybrid architectures that integrate other modules. Bello’s work [2] introduces approximate content attention with a positional attention component. Child et al. [69] observe that many early layers in the network learn locally connected patterns akin to convolutions, indicating that hybrid architectures inspired by both transformers and convolutional networks are a compelling design choice. Several recent studies explore this approach for various tasks [59, 70, 65, 25]. In contrast to prior work, we do not introduce a new attention structure. Instead, we offer theoretical proof for optimizing existing attention mechanisms. This resulting optimization approach is applicable across various attention designs and demonstrates strong performance across multiple datasets.
@@ 3 Preliminaries
SAM. Segment Anything Model (SAM) [38] consists of three core modules: image encoder, prompt encoder, and mask decoder. The SAM model has been trained on the extensive SA-1B dataset [38], which comprises more than 1 billion automatically generated masks. This dataset contains 400 times more masks than any existing segmentation dataset (COCO [45], LVIS [27] and ADE20K [82]) and is accompanied by 11 million images and 1 billion masks. Consequently, SAM exhibits valuable and robust zero-shot generalization to new data without necessitating further training. For further details regarding the SAM methodology, readers are encouraged to refer to [38]. Our Scalable Bias-Mode Attention Mask (BA-SAM) focuses its optimization on the image encoder while keeping the structures of the mask decoder and prompt encoder unchanged.
Scaling Factor in Attention. The two most commonly used attention functions are additive attention [1] and dot-product attention [62]. The vanilla Transformer chooses dot-product attention because it is much faster and more space-efficient in practice. However, for large values of the dimension of the tokens $d_{k}$, the dot products grow large in magnitude, pushing the softmax function into regions with minimal gradients. They use the scaling factor $\lambda_{d}=\frac{1}{\sqrt{d_{k}}}$ to scale the dot products. To better analyze the role of the scaling factor, we express the output element $\mathbf{O}_{i}$ and the weight coefficient $a_{i,j}$ as follows:
$\mathbf{O}_{i}$ $=\sum_{j=1}^{N}a_{i,j}x_{j}v_{j},\quad a_{i,j}=\frac{e^{\lambda_{d}q_{i}\cdot k_{j}}}{\sum_{j=1}^{N}e^{\lambda_{d}q_{i}\cdot k_{j}}},$ (1)
@@
Absolute & Relative Position Encoding in Attention. The original Transformer [62] incorporates absolute non-parametric positional encoding $p=(p_{1},\dots,p_{n})$ with the input token embedding $x$ as $x_{i}=x_{i}+p_{i}$, where $p_{i},x_{i}\in\mathrm{R^{d_{x}}}$. Other works replace them with parametric encoding [23] or adopted Fourier-based kernelized versions [51]. Absolute position encoding enforces a fixed size for all inputs. Recent work [56] considers the pairwise relationships between elements. Relative relation is important for tasks where the relative ordering or distance of the elements matters. This method encodes the relative position between input elements $x_{i}$ and $x_{j}$ into vectors $p_{i,j}^{v},p_{i,j}^{q},p_{i,j}^{k}\in\mathrm{R}^{d_{z}}$. Then, we reformulate Eq. (1) as follows:
$\mathbf{O}_{i}=\sum_{j=1}^{n}a_{i,j}\left(x_{j}v_{j}+p_{i,j}^{v}\right),$ (2)
@@ 4 Methodology
From the above section, it is evident that the original SAM’s design has a substantial limit to handling the length extrapolation problem. To address this, as shown in Figure 2, we present a Scalable Bias-mode Attention Mask (BA-SAM). In Sec. 4.1, we provide a theoretical explanation for the scaling factor used in the original Transformer and introduce a new scaling factor to regulate the magnitude inconsistency caused by length extrapolation. In Sec. 4.2 we design a bias-mode attention mask to place more focus on neighboring tokens, ensuring the sensitivity of the model when meeting positional encoding variations. Finally, we explain how we will embed our proposed BA-SAM into the SAM-based structure in Sec. 4.3.
@@ 4.1 New Scaling Factor
We discovered from the original attention module in SAM [62] that when the dot product becomes significantly large in magnitude, it can drive the Softmax layer into regions with minimal gradients. This is because the attention distribution after Softmax becomes highly concentrated, giving rise to the issue of vanishing gradients. Upon closer examination of Eq. (1), it is obvious that the computation of the $q\cdot k$ term is intrinsically tied to both the token sequences length $N$ and the dimension $d_{k}$. When the token sequence length $N$ and/or the dimension $d_{k}$ significantly increases, the overall efficacy of the attention is impacted, thus leading to a noticeable performance reduction.
To address this, we attempt to design a new scaling factor that allows the model to cope with significant variations in $N$ and $d_{k}$ values. When $N$ or $d_{k}$ grows significantly, we expect to regulate the magnitude of the values within the attention layer, maintaining a similar magnitude. While [62] introduced a scaling factor $\lambda=\frac{1}{\sqrt{d_{k}}}$ to counteract the effect of the large growth in magnitude due to the dot products, the authors did not explain how this scaling factor was designed in the original paper. Below we will provide a theoretical derivation of this scaling factor, and then elaborate on our proposed new scaling factor.
The dimension $d_{k}$. We assume that the components of $q$ and $k$ are independent Gaussian distribution variables with mean 0 and variance 1, then we want to prove that the $q\cdot k$ has mean 0 and variance $d_{k}$. The mean of $q\cdot k$ is:
@@
Given this, we can approximately consider the $q\cdot k$ values to be within the range of $-3\sqrt{d_{k}}$ to $3\sqrt{d_{k}}$, according to properties of Gaussian distribution. For larger models, $d_{k}$ is generally a larger positive value, resulting in a significant increase in the magnitude of numerical values of $q\cdot k$, compared to the additive attention option which has the range of $\left[-3,3\right]$. Consequently, the attention distribution after Softmax becomes highly concentrated. This leads to severe gradient vanishing, which in turn hampers the effectiveness of training and can lead to less desired performance. As the $q\cdot k$ values lie in the range of $[-3\sqrt{d_{k}},3\sqrt{d_{k}}]$, the scaling factor can be simply defined as $\lambda_{d}=\frac{1}{\sqrt{d_{k}}}$, in order to maintain a similar magnitude.
Our new scaling factor. We provided the interpretation on how the original scaling factor was designed, which has been overlooked by the original authors. Now we explain the design of our new scaling factor.
Our goal is to regulate the magnitude of the Softmax value $a_{ij}$ (Eq. (1)) as the token sequences length $N$ and the dimension $d_{k}$ change. We further denote the token sequences length during the training as $N_{train}$ and the token sequence length during the testing as $N_{test}$, where $N_{test}>>N_{train}$. We now consider how to maintain a similar magnitude of $a_{ij}$ during training and testing times. We introduce the scaling factor $\lambda_{N}$, which is related to $N$ and controls the output scale of the Softmax value $a_{ij}$ , namely, $S(a_{i,j}^{N_{test}})=S(a_{i,j}^{N_{train}})$. Here, $S()$ represents the scale estimation function.
We next substitute Eq. (1) into $S(a_{i,j}^{N_{test}})=S(a_{i,j}^{N_{train}})$, and simplify $q_{i}\cdot k_{j}$ into $x_{i,j}$ provided the magnitude of $q\cdot k$ has been already maintained. Now we have:
$S(\frac{e^{\lambda_{n}x_{i,j}}}{\sum_{j=1}^{n}e^{\lambda_{n}x_{i,j}}})=S(\frac{e^{x_{i,j}}}{\sum_{j=1}^{m}e^{x_{i,j}}})$ (6)
$\Rightarrow S({e^{\lambda_{n}}})=S(\frac{N_{test}}{N_{train}})\Rightarrow\lambda_{n}=log_{N_{train}}N_{test}$
Considering both $\lambda_{d}$ and $\lambda_{n}$, we can ultimately derive our new scaling factor as:
@@
Our new scaling factor in Eq. (7) ensures the magnitude of values during attention computation remains consistent, regardless of variations in $d_{k}$ and $N$. It will enhance the extrapolative capacity of the model.

### 19 Mar 2024 v3

@@ Abstract
Figure 1: Top: contrast between prior methods [81, 4] and BA-SAM. For large-scale datasets, previous approaches often resize images or change patch sizes to handle the issue of varying resolutions. In contrast, we propose a Scalable Bias-Mode Attention Mask (BA-SAM), which enhances SAM’s adaptability to varying image resolutions while eliminating structure modifications. Bottom (left): We introduce a generalized model that outperforms state-of-the-art methods across four datasets. Bottom (right): With resolution variations, prior models’ performance degrades drastically. Instead, BA-SAM consistently alleviates this issue (The evaluation metric is MAE).
@@ 1 Introduction
Recently, the computer vision community [29, 66, 28, 9, 75, 26, 87, 86, 88, 85, 89, 90, 91, 30, 92, 22, 58, 50, 49, 93] has experienced a surge in the development of various foundation models [54, 33, 21]. Notably, Meta has introduced SAM (Segment Anything Model) [36], a prompt model that has made a significant impact. SAM can segment any object in an image or video by incorporating a single visual prompt, such as a box or a point, without requiring additional training. SAM is trained on an extensive SA-1B dataset [36], consisting of over 11 million images and one billion masks. Its emergence has undeniably showcased robust generalization capabilities across diverse images and objects, paving the way for new possibilities and avenues in intelligent image analysis and understanding [78, 8, 81, 32]. Based on SAM, Some variants have been proposed, such as MobileSAM [78] and SAM-Adapter [8]. These efforts typically focus on improving SAM’s performance on specific datasets.
@@
Since SAM adopts the standard Vision Transformer [14] architecture, there are two common approaches to address the inconsistency between training and inference sizes for the ViT architecture. As depicted in Fig. 1, the first approach, e.g., MSA [81] and SAM-Adapter [8], involves directly resizing all datasets to match a predefined size. Conversely, the second approach, exemplified by FlexiViT [4], entails adjusting the patch size to accommodate larger image resolutions. Nevertheless, tuning the image or patch size necessitates a complete retraining of the model, which is cost-expensive and unacceptable for deployment in the downstream tasks. Besides, it prevents leveraging the rich prior knowledge reserved in the pre-trained model of SAM. As such, our objective is to explore a solution that enhances SAM’s adaptability to datasets of varying resolutions while avoiding structural modifications to SAM.
In this paper, we introduce a novel perspective that reframes the challenge of image resolution variation as a length extrapolation problem. Specifically, as depicted in Fig. 1, for images of varying sizes, we employ different token sequence lengths while keeping a consistent patch size. It has been observed that the inconsistency in token length between training and prediction is a key factor in performance degradation. This inconsistency manifests in two aspects: Firstly, changes in token length lead to variations in the magnitude of attention module values. When the dot product result becomes significantly large in magnitude, it can drive the subsequent Softmax layer into regions with minimal gradients. Consequently, the attention distribution after Softmax becomes highly concentrated, giving rise to the issue of vanishing gradients. Secondly, longer predictions rely on untrained information, such as additional position encodings. The introduction of untrained parameters brings a considerable amount of noise to the model, which, in turn, affects its performance.
@@
Our approach demonstrates efficacy in two scenarios: zero-shot and fine-tuning. Extensive evaluations on datasets from five diverse tasks are conducted, including DIS5K [53], DUTS [64], ISIC [12], COD10K [16], and COCO [45]. These datasets vary in resolutions, mostly exceeding SAM’s default resolution of 1024. In the zero-shot setting, our BA-SAM alleviates the model’s performance degradation caused by expanding the inference resolution without requiring any additional training. With a few fine-tuning epochs on downstream tasks, our BA-SAM consistently achieves state-of-the-art accuracy across all datasets. Additionally, to further demonstrate BA-SAM’s generalizability, we propose a generalized model and a new benchmark, which utilize one model to attain state-of-the-art performance across all four datasets simultaneously.
@@ 2 Related Work
Visual Foundation Models. Models that are trained on broad and can be adapted to numerous downstream tasks are called “Foundation Models” [5, 67, 43, 40, 72, 77, 84, 41]. These models, Vision-Language Models (VLM) (CLIP [54] and DALL-E [55]) combine computer vision and natural language processing to understand and generate descriptions or analyze visual content using textual and visual information. Masked Image Modeling [74, 48] (MIM) masks parts of an image during the training to encourage a model to learn contextual information and complete missing regions. SAM [36] is a model designed for segmenting objects or areas in images, offering precise segmentation capabilities. We use a variant of SAM called MobileSAM [78] as the baseline method.
@@
Figure 2: Illustration of the proposed BA-SAM method. (a) In the original SAM, when input token sequences length varies during testing, the magnitude of the Softmax outputs changes drastically. We propose a new scaling factor to address this issue. (b) We introduce a bias-mode attention mask, which produces increasing penalties on attention scores as the distance between the query and key grows.
@@
Resolution Variation Processing. To enable models to be more adaptable to variations in resolutions, traditional methods to deal with VIT have relied on adjustments to positional embeddings [38] and patch sizes [4, 6, 24, 76, 44, 29, 37]. Patch n’ Pack [13] employed sequence packing during the training to handle inputs with arbitrary resolutions and aspect ratios. All of them necessitate training from scratch, incurring substantial computational and time costs. In contrast to previous approaches, we extend the concept of length extrapolation from NLP into the context of addressing scale variations in CV. Length extrapolation refers to a model’s ability to generalize well to longer inputs than those it was trained on. In NLP, it has been successfully used, such as in ALIBI [52] and KERPLE [11], to enable models to adapt to longer sequences without significant performance degradation. Our approach seamlessly extends to two scenarios: zero-shot and finetuning, allowing us to leverage prior knowledge embedded in the SAM and significantly reduce training efforts.
@@
Parameter Efficient Tuning. There have been some pioneering works for the Parameter Efficient Tuning (PEFT) of visual models, such as AdaptFormer [7] and visual prompt tuning (VPT) [34]. He et al. [27] analyzed the unified view among PETL techniques such as prefixtuning [42], Prompt-tuning [34], and adapter [7]. Our method belongs to the category of Parameter Efficient Tuning.
@@
Visual Attention Modeling. Various studies have incorporated attention mechanisms into neural network architectures designed for visual tasks [63, 68, 39, 3, 31, 96]. These mechanisms are employed in a channel-wise manner to capture cross-feature information [66, 10, 82]. They are also used for selecting paths in different branches of a network [61], or a combination of both strategies [80]. The advent of transformers has led to hybrid architectures that integrate other modules. Bello’s work [2] introduces approximate content attention with a positional attention component. Child et al. [69] observe that many early layers in the network learn locally connected patterns akin to convolutions, indicating that hybrid architectures inspired by both transformers and convolutional networks are a compelling design choice. Several recent studies explore this approach for various tasks [59, 70, 65, 25]. In contrast to prior work, we do not introduce a new attention structure. Instead, we offer theoretical proof for optimizing existing attention mechanisms. This resulting optimization approach is applicable across various attention designs and demonstrates strong performance across multiple datasets.
@@ 3 Preliminaries
SAM. Segment Anything Model (SAM) [36] consists of three core modules: image encoder, prompt encoder, and mask decoder. It has been trained on SA-1B dataset [36], which comprises more than 1 billion automatically generated masks. Consequently, SAM exhibits valuable and robust zero-shot generalization to new data without necessitating further training. further details can refer to [36]. Our Scalable Bias-Mode Attention Mask (BA-SAM) focuses its optimization on the image encoder while keeping the structures of the mask decoder and prompt encoder unchanged.
Attention in Transformer. In this work, we define the input sequence of image patches, $\mathbf{x}=(\mathbf{x}_{1},\dots,\mathbf{x}_{n})$ with length $N$, where $\mathbf{x}_{i}\in\mathbb{R}^{d_{x}}$. $q_{i}$, $k_{j}$,$v_{j}$ are calculated by $\mathbf{x}_{i}\mathbf{W}^{Q}$ , $\mathbf{x}_{j}\mathbf{W}^{K}$, $\mathbf{x}_{j}\mathbf{W}^{V}$. Here, the projections $\mathbf{W}^{Q},\mathbf{W}^{K},\mathbf{W}^{V}\in\mathbb{R}^{d_{x}\times d_{k}}$ are parameter matrices.
(i) Scaling Factor. The two most commonly used attention functions are additive attention [1] and dot-product attention [62]. The vanilla Transformer chooses dot-product attention for its space efficiency in practice. However, for larger values of $d_{k}$, the dot products grow large in magnitude, pushing the Softmax function into regions with minimal gradients. They use scaling factor $\lambda_{d}=\frac{1}{\sqrt{d_{k}}}$ to scale the dot products, where $d_{k}$ denotes the dimension. To better analyze the role of scaling factor, we express the output element $\mathbf{O}_{i}$ and the weight coefficient $a_{i,j}$ as follows:
$\mathbf{O}_{i}$ $=\sum_{j=1}^{N}a_{i,j}v_{j},\quad a_{i,j}=\frac{e^{\lambda_{d}q_{i}\cdot k_{j}}}{\sum_{j=1}^{N}e^{\lambda_{d}q_{i}\cdot k_{j}}},$ (1)
@@
(ii) Absolute & Relative Position Encoding. The original Transformer [62] incorporates absolute non-parametric positional encoding $p=(p_{1},\dots,p_{n})$ with $x$ as $x_{i}=x_{i}+p_{i}$. Other works replace them with parametric encoding [23] or adopted Fourier-based kernelized versions [51]. Absolute position encoding enforces a fixed size for inputs. Recent work  [56] considers the pairwise relationships between elements, which encodes the relative position between input $x_{i}$ and $x_{j}$ into vectors $p_{i,j}^{v},p_{i,j}^{q},p_{i,j}^{k}\in\mathrm{R}^{d_{k}}$. Then, we reformulate Eq. (1) as follows:
$\mathbf{O}_{i}=\sum_{j=1}^{N}a_{i,j}\left(v_{j}+p_{i,j}^{v}\right),$ (2)
@@ 4 Methodology
Based on the preliminaries, we further analyze the characteristics of SAM: the original SAM sets the input to a fixed resolution of 1024, where it uses absolute position coding and the dot product. Thus, there are significant limitations in the processing of length extrapolation problems. To address this, as shown in Fig. 2, we present a Scalable Bias-mode Attention Mask (BA-SAM). In Sec. 4.1, we provide a theoretical explanation for the scaling factor used in the original Transformer and introduce a new scaling factor to regulate the magnitude inconsistency caused by length extrapolation. In Sec. 4.2 we design a bias-mode attention mask to place more focus on neighboring tokens, mitigating the impact of untrained distant information. Finally, we explain how we will embed our BA-SAM into the SAM-based structure in Sec. 4.3.
@@ 4.1 New Scaling Factor
We observe that in the original attention module of SAM [62] when the dot product becomes significantly large in magnitude, it can drive the Softmax layer into regions with minimal gradients. This is because the attention distribution after Softmax becomes highly concentrated, giving rise to the issue of vanishing gradients. Upon closer examination of Eq. (1), it is obvious that the computation of the $q\cdot k$ the term is intrinsically tied to both the token sequences length $N$ and the dimension $d_{k}$. When the token sequence length $N$ and the dimension $d_{k}$ significantly increase, the overall efficacy of the attention is affected, thus leading to a noticeable performance degradation.
To address this issue, we attempt to design a new scaling factor that allows the model to cope with variations in $N$ and $d_{k}$. When $N$ or $d_{k}$ grows significantly, we expect to regulate the magnitude of the values within the attention layer, maintaining a similar magnitude. [62] introduced a scaling factor $\lambda=\frac{1}{\sqrt{d_{k}}}$ to counteract the effect of the large growth in magnitude due to the dot products. Below we will provide a theoretical derivation of this scaling factor, and then elaborate on our proposed new scaling factor.
The dimension $d_{k}$. Following work [62], we assume that the components of $q$ and $k$ are independent random variables with mean 0 and variance 1. The mean of $q\cdot k$ is:
@@
Given this, we can approximately consider the $q\cdot k$ values to be within the range of $-3\sqrt{d_{k}}$ to $3\sqrt{d_{k}}$, according to properties of Gaussian distribution. For larger models, $d_{k}$ is generally a larger positive value, resulting in a significant increase in the magnitude of numerical values of $q\cdot k$, compared to the additive attention option, which has the range of $\left[-3,3\right]$. Consequently, the attention distribution after Softmax becomes highly concentrated. This leads to severe gradient vanishing, which hampers the effectiveness of the training and can lead to less desired performance. As the $q\cdot k$ values lie in the range of $[-3\sqrt{d_{k}},3\sqrt{d_{k}}]$, the scaling factor can be simply defined as $\lambda_{d}=\frac{1}{\sqrt{d_{k}}}$, in order to maintain a similar magnitude.
Our new scaling factor. We have provided the interpretation of how the original scaling factor was designed. Now, we explain the design of our new scaling factor.
According to Eq. 4 and Eq. 5, the scale of $q\cdot k$ has been consistent with the additive attention by $\lambda_{d}$, which can be seen as $a_{i,j}$ is independent of $d_{k}$. We simplify $\lambda_{d}q_{i}\cdot k_{j}$ into $x_{i,j}$ and further discuss the effect of length $N$ on $a_{i,j}$.
In Eq.1 , $a_{i,j}$ can be seen as the conditional distribution with $i$ as the condition and $j$ as the random variable. Inspired by [60], we introduce information entropy to constrain $a_{i,j}$. Specifically, entropy is a measure of uncertainty, and we expect the uncertainty of $a_{i,j}$ to be insensitive to the length $N$ (i.e., the value of each $a_{i,j}$ will change when the token increases, but the entropy value of the overall $a_{i,j}$ can remain relatively stable). The entropy of $a_{i,j}$ is :$\mathcal{H}_{i}=-\sum_{j=1}^{N}a_{i,j}\log a_{i,j}$ and we substitute Eq. (1):
$\mathcal{H}_{i}=\log\sum_{j=1}^{N}e^{\lambda_{n}x_{i,j}}-\frac{\sum_{j=1}^{N}e^{\lambda_{n}x_{i,j}}\left(\lambda_{n}x_{i,j}\right)}{\sum_{j=1}^{N}e^{\lambda_{n}x_{i,j}}}$ (6)
Then, we substitute the approximate estimates into Eq. (6):
$\sum_{j=1}^{N}e^{x_{i,j}}=N\times\frac{1}{N}\sum_{j=1}^{N}e^{x_{i,j}}\approx N\mathbb{E}_{j}\left[e^{x_{i,j}}\right]$ (7)
$\mathbb{E}_{j}\left[e^{x_{i,j}}\left(x_{i,j}\right)\right]\approx 0,\mathbb{E}_{j}\left[e^{x_{i,j}}\right]=O(1)$
We wish to use $\lambda_{n}$ to offset the effect of $N$ on $\mathcal{H}_{i}$. Then, we have the following result:
$\mathcal{H}_{i}\approx\log N-k\lambda_{n}=0\Rightarrow\lambda_{n}=\frac{\log N}{k}$ (8)
where $k$ is a parameter value. We denote the token sequence length during the training as $N_{train}$ and the token sequence length during the testing as $N_{test}$, where $N_{test}>>N_{train}$. When $N=N_{train}$, $\lambda_{n}=1$ (consistent with the training length). As such, $k=\log N_{train}$ and finally we have $\lambda_{n}=log_{N_{train}}N_{test}$. Considering both $\lambda_{d}$ and $\lambda_{n}$, we can ultimately derive our new scaling factor as:
@@
Our new scaling factor in Eq. (9) ensures attention computation remains consistent, regardless of variations in $d_{k}$ and $N$. It will enhance the extrapolative capacity of the model.